Let {an} be an arithmetic sequence such that its 1st, 20th, and 58th terms are consecutive terms of some geometric sequence. Find the common ratio of the geometric sequence.


1
Expert's answer
2021-08-30T16:21:27-0400

Solution:

"a_1,a_{20},a_{58}" are in GP.

Let the first term "a_1" be "a" and the common difference is "d".

"a_{20}=a+19d
\\a_{58}=a+57d"

When p,q,r are in GP, then "q^2=pr"

So, "a_{20}^2=a_1\times a_{58}"

"\Rightarrow (a+19d)^2=a(a+57d)
\\ \Rightarrow a^2+361d^2+38ad=a^2+57ad
\\ \Rightarrow 361d^2-19ad=0
\\\Rightarrow 19d(19d-a)=0
\\\Rightarrow d=0;d=\dfrac{a}{19}"

Rejecting d=0 as it will give trivial AP.

So, "a_1=a,a_{20}=a+19(\dfrac{a}{19})=2a,a_{58}=a+57(\dfrac{a}{19})=4a"

Now, a, 2a, 4a are in GP.

Common ratio"=r=\dfrac{2a}{a}=2"